Find the total derivative of Z with respect to x. z = x²y - y2x - 5x3, where y = (3x + 5)12
Answer #1

12 Given - z= x²y - y2x - 5x3 ;y= (3x+5) ² > Z = x2 (3x + 5)2 - *(3x+5)24 - 5x3 = 22•(3x+5)2 + x2. 12 (32+5).3 - (3x+5) 24 –di = (3x+5) (6x2 +10% + 36x2) - - (3x+5) 23 (3x+ 5 + 72%) - 15x2 (3x+5) (42x2 +10%) - (22+5) 23 (75x+5) -15x2 - 2x (21 x+5)